Give the equation of the horizontal asymptote of the rational function.
g(x) = (3)/(x+4)
A) y = 3
B) y = 0
C) y = 3/4
D) x = -4

Answer by stanbon(75887)

Give the equation of the horizontal asymptote of the rational function.
g(x) = (3)/(x+4)
Horizontal asymptote : y = 0*x/1*x = 0/1 = 0
---
Another method:
Divide 0 by (x+4).
You get a quotient of 0 and a remainder of 3/(x+4)
y = the quotient is the horizontal asymptote
Ans: y = 0
=======================================
Cheers,
Stan H.
=======================================
x = 0
A) y = 3
B) y = 0
C) y = 3/4
D) x = -4